OpenStudy (learner31):

use the sum to product identities to rewrite cos15+cos19 as a product

OpenStudy (holsteremission):

Well, you probably know the identity \[\cos x+\cos y=2\cos\frac{x+y}{2}\cos\frac{x-y}{2}\]So what's the problem?